The charter package^{∗}

Peter Dyballa

1994/04/07

1 Introduction

This file contains the external font information needed to load the Charter font, and a suitable package file to utilize it.

From this file all .fd files (font definition files) for the PostScript fonts with old encoding (OT1) are created, as well as those with the new encoding (T1).

2 Customization

The usual customization will be to add support for additional combinations in this file, such as a derived bold slanted font.

∗This file has version number 2, last revised 1994/04/07.